# Research impact evaluation based on effective authorship contribution sensitivity: h-leadership index

### v4

@@ Abstract
[…]
[inst3]organization=Centre for Artificial Intelligence and Innovation,addressline=Pingla Institute, city=Sydney, country=Australia
@@ 4.1 Data and bibliometrics
Table 2 presents the hl-index for leading researchers defined by citations received on researchers’ Google Scholar profiles in the selected Go8 Australian universities. For instance, Google Scholar provides a list of the top 10 researchers at UniMelb based on citations with links on the page to the next 10 researchers  [^5] https://scholar.google.com/citations?view_op=view_org&hl=en&org=10194437184893062620. The reason for using Google Scholar for top author identification instead of relying on Scopus has been explained in Section 3.2. We selected 10 out of the top 50 profiles in this study based on random sampling from each of the Go8 Australian universities in Table 2. We note that the top-cited researchers at Australian National University (ANU) and Monash University engage in Physics, Astronomy and Medicine disciplines (Figure 9), which explains the higher number of co-authors observed. We also note that Scopus search API clips a publication’s authors to the first 100, and obtaining a full list requires using abstract retrieval API [^6] https://cran.r-project.org/web/packages/rscopus/vignettes/multi_author.html. Since an abstract retrieval API call needs to be made per publication, we decided against its usage due to existing weekly quota limits. We highlight that in our selected random samples, we have 2 author profiles with 100 median coauthors at ANU and 3 at Monash. We notice huge variations when comparing the h-index to our hl-index (59 vs 4) for the 3rd ANU profile and the 5th ANU profile (80 vs 5). We also notice similar patterns when comparing the h-index and hl-index for Monash profiles (1st, 3rd, and 5th profiles. We provide a complete list of the top 50 profiles in metrics.csv for all Go8 universities via our GitHub repository  [^7] https://github.com/nepython/metrics/tree/main/results. We also notice that the first, single and last authorship makes a large impact on the hl-index when compared to the h-index, since authorship position is the key principle of the hl-index. In cases when a medium number of coauthors is low (2 or 3), then the hl-index is close to the h-index.
@@
Median citations h-index hl-index % first, single, last authorship Median author position Median co-authors University
19 32 26 27.10 4 9 UniMelb
27 57 39 30.87 7 10 UniMelb
11 38 32 38.59 3 6 UniMelb
43 126 126 58.88 3 5 UniMelb
23 104 104 62.97 3 4 UniMelb
36 83 83 44.80 5 7 UniMelb
36 101 101 71.80 3 4 UniMelb
30 116 116 85.75 2 4 UniMelb
51 134 134 76.68 5 6 UniMelb
24 64 64 46.54 2 4 UniMelb
7 61 61 91.38 3 3 USyd
25 99 99 57.65 3 4 USyd
8 26 20 34.05 7 18 USyd
13 12 12 36.36 3 7 USyd
38 120 116 29.62 4 7 USyd
14 103 102 41.59 3 6 USyd
6 79 79 83.04 4 4 USyd
25 122 121 27.65 6 15 USyd
15 62 59 32.71 4 7 USyd
26 101 101 55.07 6 7 USyd
36 29 28 40.00 2 5 ANU
12 135 135 81.32 4 4 ANU
65 59 4 0.89 4 100 ANU
52 78 75 25.21 13 21 ANU
72 80 5 4.00 2 100 ANU
28 87 87 73.15 3 3 ANU
26 78 77 60.54 3 5 ANU
18 25 19 26.56 3 8 ANU
20 62 62 90.53 2 3 ANU
40 65 64 49.26 4 6 ANU
35 68 9 5.29 4 100 Monash
13 79 78 41.93 5 7 Monash
26 116 15 2.24 4 100 Monash
26 41 41 65.81 2 4 Monash
36 49 37 20.51 37 100 Monash
10 89 87 55.41 4 8 Monash
17 103 102 40.31 5 7 Monash
13 12 12 43.48 2 5 Monash
50 89 62 29.52 4 12 Monash
25 104 87 11.55 12 25 Monash
21 65 63 97.45 2 3 UNSW
21 86 85 23.76 6 9 UNSW
28 68 67 45.66 2 4 UNSW
20 55 55 50.61 4 6 UNSW
16 92 92 63.86 5 6 UNSW
47 34 32 46.44 2 5 UNSW
18 127 127 98.64 3 4 UNSW
24 114 114 69.12 3 6 UNSW
16 84 78 29.43 7 10 UNSW
26 45 45 25.62 4 8 UNSW
@@
Table 2: h-leadership index for 10 randomly sampled researchers across a) UniMelb, b) USyd, c) ANU, d) Monash and e) UNSW. We have not listed any metrics such as citations or number of publications to protect author privacy. We derive the median citation number by sorting an author’s publications by citation and considering the median publication’s citation count. We derive the median author position by sorting publications based on authorship position. We sorted publications by the number of coauthors to obtain the median number of coauthors.

### v5

@@ 4.1 Data and bibliometrics
Table 2 presents the hl-index for leading Medicine researchers defined by citations received on researchers’ Google Scholar profiles in the selected Go8 Australian universities. For instance, Google Scholar provides a list of the top 10 researchers at UniMelb based on citations with links on the page to the next 10 researchers  [^5] https://scholar.google.com/citations?view_op=view_org&hl=en&org=10194437184893062620. The reason for using Google Scholar for top author identification instead of relying on Scopus has been explained in Section 3.2. Similarly, results for Engineering discipline are shown in Table 3 and data for all disciplines available in code repository. Discipline specific analysis is important as citation practices vary significantly between disciplines. We note that the top-cited researchers at Australian National University (ANU) and Monash University engage in Physics, Astronomy and Medicine disciplines (Figure 9), which explains the higher number of co-authors observed. We also note that Scopus search API clips a publication’s authors to the first 100, and obtaining a full list requires using abstract retrieval API [^6] https://cran.r-project.org/web/packages/rscopus/vignettes/multi_author.html. Since an abstract retrieval API call needs to be made per publication, we decided against its usage due to existing weekly quota limits. We highlight that in our selected random samples, we have 2 author profiles with 100 median coauthors at ANU and 3 at Monash. We notice huge variations when comparing the h-index to our hl-index (59 vs 4) for the 3rd ANU profile and the 5th ANU profile (80 vs 5). We also notice similar patterns when comparing the h-index and hl-index for Monash profiles (1st, 3rd, and 5th profiles. We provide a complete list of the top 50 profiles in metrics.csv for all Go8 universities via our GitHub repository  [^7] https://github.com/nepython/metrics/tree/main/results. We also notice that the first, single and last authorship makes a large impact on the hl-index when compared to the h-index, since authorship position is the key principle of the hl-index. In cases when a medium number of coauthors is low (2 or 3), then the hl-index is close to the h-index.
@@
Name Median citations h-index hl-index i10-index hm-index % first, single, last authorship Median number of co-authors University
Jenkins, Mark A. 26 84 79 352 19 15.83 19 UniMelb
Loi, Sherene M. 36 102 101 238 19 36.18 15 UniMelb
Herrman, Helen Edith 12 49 49 157 19 46.69 5 UniMelb
Wei, Andrew H. 13 55 55 139 12 38.46 13 UniMelb
Seymour, John Francis 18 104 102 364 28 39.23 8 UniMelb
Macaskill, Petra 34 66 66 145 20 15.85 6 USyd
Halliday, Glenda Margaret 33 117 114 559 47 34.46 7 USyd
Wang, Jiejin Jie J. 38 120 116 586 44 28.45 7 USyd
Bauman, Adrian Ernest 21 121 121 735 54 40.49 6 USyd
Young, Jane M. 20 52 51 174 23 31.58 5 USyd
Robinson, Bruce W.S. 27 63 63 197 24 50.58 5 UWA
Prince, Richard L. 23 70 68 240 25 56.90 7 UWA
Liu, Jianjun 34 106 87 385 16 14.53 20 UWA
Jablensky, Assen Verniaminov 19 73 65 213 30 57.86 5 UWA
Hankey, Graeme J. 22 150 119 614 47 35.61 7 UWA
Fitzgerald, Paul B. 27 91 91 428 40 54.34 6 ANU
Griffiths, Kathleen Margaret 40 72 72 171 29 52.20 4 ANU
Mann, Graham J. 28 89 61 257 16 13.91 16 ANU
Colquhoun, Samantha M. 28 28 23 38 4 12.50 9 ANU
Nolan, Christopher James 17 37 36 66 13 35.65 6 ANU
Norman, Robert John 22 102 101 359 41 54.17 5 UoA
Price, Timothy J. 10 54 53 170 15 36.36 9 UoA
Amare, Azmeraw T. 55 58 46 88 4 13.56 100 UoA
Munn, Zachary 9 42 40 100 15 52.71 6 UoA
Sanders, Prashanthan 14 104 103 433 29 39.34 8 UoA
Lalloo, Ratilal 13 52 27 106 12 39.78 5 UQ
Evans, David M. 47 103 86 274 18 20.88 21 UQ
Roberts, Jason A. 15 87 86 411 33 47.38 7 UQ
Montgomery, Grant W. 39 158 120 673 30 16.14 16 UQ
Charlson, Fiona J. 240 55 40 68 5 26.74 19 UQ
Cumpston, Miranda S. 13 12 12 14 3 43.48 5 Monash
Bailey, Michael J. 19 95 95 444 33 12.21 7 Monash
Li, Shanshan 25 74 57 255 16 16.53 12 Monash
McNeil, John J. 13 79 78 398 31 39.85 7 Monash
Cheng, Allen C. 15 74 73 312 27 34.42 7 Monash
Mattick, Richard P. 21 62 62 198 25 49.64 6 UNSW
Carr, Vaughan James 24 72 67 226 25 38.60 7 UNSW
Law, Matthew G. 21 86 85 395 27 22.36 9 UNSW
Brodaty, Henry 20 105 101 574 42 40.67 7 UNSW
Anstey, Kaarin J. 23 83 83 391 41 54.53 5 UNSW
@@
Table 2: h-leadership index for 5 randomly sampled top researchers in ”Medicine” discipline at each Go8 institute. We derived the median citation number by sorting an author’s publications by citation and considering the median publication’s citation count. We sorted publications by the number of coauthors to obtain the median number of coauthors.
Name Median citations h-index hl-index i10-index hm-index % first, single, last authorship Median number of co-authors University
Yabsley, Bruce D. 34 132 3 1109 10 0.21 100 USyd
Eggleton, Benjamin J. 1 97 96 477 42 66.98 6 USyd
Driscoll, Tim Robert 22 72 48 157 17 30.00 7 USyd
Mai, Yiu Wing 20 109 109 732 61 69.65 3 USyd
Zomaya, Albert Y.H. 6 79 79 487 40 75.09 4 USyd
Colless, Matthew M. 46 89 87 224 17 12.68 18 ANU
Zheng, Liang 31 51 51 85 18 45.59 4 ANU
Hartley, Richard I. 20 62 62 173 39 64.02 3 ANU
Yin, Zongyou 28 70 70 151 17 38.22 8 ANU
McClelland, David E. 28 105 36 355 14 21.05 100 ANU
Hessel, Volker 14 77 77 346 36 49.01 5 UoA
Qiao, Shizhang Zhang 84 175 175 492 58 71.56 6 UoA
Govindan, Kannan 39 110 110 326 55 49.53 4 UoA
Jackson, Paul D. 44 139 12 1199 12 0.89 100 UoA
Jiao, Yan 72 69 69 113 16 17.65 6 UoA
Milburn, G. J. 15 78 77 238 44 76.18 3 UQ
Yamauchi, Yusuke 33 147 147 945 59 51.83 8 UQ
Keller, Jurg 59 104 104 262 41 44.87 5 UQ
Sawadsky, Andreas 84 57 5 78 1 2.27 100 UQ
Chen, Guanrong (Ron) 15 146 146 1021 82 71.35 3 UQ
Nash, Jordan A. 33 141 27 1590 14 0.21 100 Monash
Skands, Peter Zeiler 20 38 31 63 14 64.76 4 Monash
Maier, Stefan A. 17 103 102 412 42 32.78 7 Monash
Egede, Ulrik 26 116 15 798 8 1.46 100 Monash
Valencia, German 15 38 34 112 22 85.71 2 Monash
Fane, A. G.(Tony) 44 119 119 471 62 55.59 4 UNSW
Ng, Derrick Wing Kwan 15 79 79 326 34 34.11 5 UNSW
Dai, Liming 19 166 166 733 69 48.90 5 UNSW
Boufous, Soufiane 25 46 33 98 12 30.66 8 UNSW
Boyer, Cyrille Andre 46 99 98 346 43 52.44 5 UNSW
Table 3: h-leadership index for 5 randomly sampled top researchers in ”Engineering” discipline at six Go8 institutions. We derived the median citation number by sorting an author’s publications by citation and considering the median publication’s citation count. We sorted publications by the number of coauthors to obtain the median number of coauthors.